# Discrete Neural Algorithmic Reasoning

### v2

@@ Abstract
Neural algorithmic reasoning aims to capture computations with neural networks via learning the models to imitate the execution of classic algorithms. While common architectures are expressive enough to contain the correct model in the weights space, current neural reasoners are struggling to generalize well on out-of-distribution data. On the other hand, classic computations are not affected by distributional shifts as they can be described as transitions between discrete computational states. In this work, we propose to force neural reasoners to maintain the execution trajectory as a combination of finite predefined states. To achieve that, we separate discrete and continuous data flows and describe the interaction between them. Trained with supervision on the algorithm’s state transitions, such models are able to perfectly align with the original algorithm. To show this, we evaluate our approach on multiple algorithmic problems and get perfect test scores both in single-task and multitask setups. Moreover, the proposed architectural choice allows us to prove the correctness of the learned algorithms for any test data.
@@ 1 Introduction
Learning to capture algorithmic dependencies in data and to perform algorithmic-like computations with neural networks are core problems in machine learning, studied for a long time using various approaches (Roni Khardon, 1994; Graves et al., 2014; Zaremba & Sutskever, 2014; Reed & De Freitas, 2015; Kaiser & Sutskever, 2015; Veličković et al., 2020b).
Neural algorithmic reasoning (Veličković & Blundell, 2021) is a research area focusing on building models capable of executing classic algorithms. Relying on strong theoretical guarantees of algorithms to work correctly on any input of any size and distribution, this setting provides unlimited challenges for out-of-distribution generalization of neural networks. Prior work explored this setup using the CLRS-30 benchmark (Veličković et al., 2022), which covers classic algorithms from the Introduction to Algorithms textbook (Cormen et al., 2009) and uses graphs as a universal tool to encode data of various types. Importantly, CLRS-30 also provides the decomposition of classic algorithms into subroutines and simple transitions between consecutive execution steps, called hints, which can be used during training in various forms.
The core idea of the CLRS-30 benchmark is to understand how neural reasoners generalize well beyond the training distribution, namely on larger graphs. Classic algorithms possess strong generalization due to the guarantee that correct execution steps never encounter ‘out-of-distribution’ states, as all state transitions are predefined by the algorithm. In contrast, when encountering inputs from distributions that significantly differ from the train data, neural networks are usually not capable of robustly maintaining internal calculations in the desired domain. Consequently, due to the complex and diverse nature of all possible data that neural reasoners can be tested on, the generalization performance of such models can vary depending on particular test distribution (Mahdavi et al., 2023).
Given that, it is becoming important to interpret internal computations of neural reasoners to find errors or to prove the correctness of the learned algorithms (Georgiev et al., 2021).
Interpretation methods have been actively developing recently due to various real-world applications of neural networks and the need to debug and maintain systems based on them. Especially, the Transformer architecture (Vaswani et al., 2017) demonstrates state-of-the-art performance in natural language processing and other modalities, representing a field for the development of interpretability methods (Elhage et al., 2021; Weiss et al., 2021; Zhou et al., 2024; Lindner et al., 2024). Based on active research on a computational model behind the transformer architecture, recent works propose a way to learn models that are fully interpretable by design (Friedman et al., 2023).
@@
In this paper, we propose to force neural reasoners to follow the execution trajectory as a combination of finite predefined states, which is important for both generalization ability and interpretability of neural reasoners. To achieve that, we start with an attention-based neural network and describe three building blocks to enhance its generalization abilities: feature discretization, hard attention and separating discrete and continuous data flows. In short, all mentioned blocks are connected:
@@
  • Hard attention is needed to ensure that attention weights will not be annealed for larger graphs. Also, hard attention limits the set of possible messages that each node can receive;
@@
Then, we build fully discrete neural reasoners for different algorithmic tasks and demonstrate their ability to perfectly mimic ground-truth algorithm execution. As a result, we achieve perfect test scores on the multiple algorithmic tasks with guarantees of correctness on any test data. Moreover, we demonstrate that a single network is capable of executing all covered algorithms in a multitask manner with perfect generalization too.
In summary, we consider the proposed blocks as a crucial component for robust and interpretable neural reasoners and demonstrate that trained with hint supervision, discretized models perfectly capture the dynamic of the underlying algorithms and do not suffer from distributional shifts.
@@ 2.1 Algorithmic reasoning
Performing algorithmic-like computations usually requires the execution of sequential steps and the number of such steps depends on the input size. To imitate such computations, neural networks are expected to be based on some form of recurrent unit, which can be applied to a particular problem instance several times (Kaiser & Sutskever, 2015; Zaremba & Sutskever, 2014; Vinyals et al., 2015; Veličković et al., 2020b).
@@
The CLRS-30 benchmark covers different algorithms over various domains (arrays, strings, graphs) and formulates them as algorithms over graphs. Also, CLRS-30 proposes to utilize the decomposition of the algorithmic trajectory execution into simple logical steps, called hints. Using this decomposition is expected to better align the model to desired computations and prevent it from utilizing hidden non-generalizable dependencies of a particular train set. Prior work demonstrates a wide variety of additional inductive biases for models towards generalizing computations, including different forms of hint usage (Veličković et al., 2022; Bevilacqua et al., 2023), biases from standard data structures (Jürß et al., 2024; Jain et al., 2023), knowledge transfer and multitasking (Xhonneux et al., 2021; Ibarz et al., 2022; Numeroso et al., 2023), etc. Also, recent studies demonstrate several benefits of learning neural reasoners end-to-end without any hints at all (Mahdavi et al., 2023; Rodionov & Prokhorenkova, 2023).
The recently proposed SALSA-CLRS benchmark (Minder et al., 2023) enables a more thorough OOD evaluation compared to CLRS-30 with increased test sizes (up to 100-fold train-to-test scaling, compared to 4-fold for CLRS-30) and diverse test distributions. Despite significant gains in the performance of neural reasoners in recent work, current models still struggle to generalize to out-of-distribution (OOD) test data (Mahdavi et al., 2023; Georgiev et al., 2023; Minder et al., 2023). While de Luca & Fountoulakis (2024) prove by construction the ability of the transformer-based neural reasoners to perfectly simulate graph algorithms (with minor limitations occurring from the finite precision), it is still unclear if generalizable and interpretable models can be obtained via learning. Importantly, the issues of OOD generalization are induced not only by the challenges of capturing the algorithmic dependencies in the data but also by the need to carefully operate with continuous inputs. For example, investigating the simplest scenario of learning to emulate the addition of real numbers, Klindt (2023) demonstrates the failure of some models to exactly imitate the desired computations due to the nature of gradient-based optimization. This limitation can significantly affect the performance of neural reasoners on adversarial examples and larger input instances when small errors can be accumulated.
@@ 2.2 Transformer interpretability and computation model
RASP (Weiss et al., 2021) is a programming language proposed as a high-level formalization of the computational model behind transformers. The main primitives of RASP are elementwise sequence functions, $select$ and $aggregate$ operations, which conceptually relate to computations performed by different blocks of the model. Later, Lindner et al. (2024) presented Tracr, a compiler for converting RASP programs to the weights of the transformer model, which can be useful for evaluating interpretability methods.
While RASP might have limited expressibility, it supports arbitrary complex continuous functions which in theory can be represented by transformer architecture, but are difficult to learn. Also, RASP is designed to formalize computations over sequences of fixed length. Motivated by that, Zhou et al. (2024) proposed RASP-L, a restricted version of RASP, which aims to formalize the computations that are easy to learn with transformers in a size-generalized way. The authors also conjecture that the length-generalization of transformers on algorithmic problems is related to the ‘simplicity’ of solving these problems in RASP-L language.
@@ 3.1 Encode-process-decode paradigm
If the model is trained with hint supervision, the changes of node and edge features at each step are expected to be related to the original algorithm execution. In this sense, the processor network is aimed to mimic the algorithm’s execution in the latent space.
@@ 3.2 Discrete neural algorithmic reasoners
Each node feature vector $x_{i}$ is projected into $query$ $(Q_{i})$, $key$ $(K_{i})$, and $value$ $(V_{i})$ vectors via learnable parameter matrices $W_{Q}$, $W_{K}$, and $W_{V}$, respectively. Edge features $e_{ij}$ are projected into $key$ $(K_{ij})$ vector with a matrix $W_{K}^{E}$. Then, for each directed edge from node $j$ to node $i$ in the graph $G$, we compute the attention coefficient
@@
where $\langle a,b\rangle$ denotes the dot product. Then, each node $i$ normalizes all attention coefficients across its neighbors with the softmax function and temperature $\tau$ and receives the aggregated message:
$\hat{\alpha}_{ij}=\frac{\exp(\alpha_{ij}/\tau)}{\sum_{k\in\textit{N}(i)}\exp(\alpha_{ik}/\tau)},\,\,\,\,\,\,M_{i}=\sum_{k\in\mathcal{N}(i)}\hat{\alpha}_{ik}V_{k},$ (1)
where $\mathcal{N}(i)$ denotes the set of all incoming neighbors of node $i$ and $M_{i}$ is the message sent to the $i$-th node.
@@
Similar to Transformer Programs, we enforce attention to be hard attention. We found this property important not only for interpretability but also for size generalization, as hard attention allows us to overcome the annealing of the attention weights for arbitrarily large graphs and strictly limits the set of messages that each vertex can receive.
After message computation, node and edge features are updated depending on the current values and sent messages using feed-forward MLP blocks:
@@
We also enforce all node and edge features to be from a fixed finite set, which we call states. We ensure such property by adding discrete bottlenecks at the end of the processor block:
@@
We implement discretization by projecting the features to the vectors of size $k$ which we force to be one-hot using annealing Gumbel-Softmax (Jang et al., 2017) during training and the argmax at the inference.
@@ 3.3 Continuous inputs
Clearly, most of the algorithmic problems operate with continuous or unbounded inputs (e.g., weights on edges). Usually, all input data is encoded into node and edge features and the processor operates over the resulting vectors. The proposed discretization of such vectors would lead to the loss of information necessary for performing correct execution steps. One possible option to operate with such inputs (we will call them scalars, meaning both continuous or size-dependent integer inputs, such as node indexes) is Neural Execution Engines (Yan et al., 2020), which allows one to operate with bit-wise representations of integer and (in theory) real numbers. Such representations are bounded by design, but fully discrete and interpretable.
We propose another option: to maintain scalar inputs (denote them by $S$) separately from the node and edge features and use them only as edge priorities $s_{ij}$ in the attention block. If scalars are related to the nodes, we assign them to edges depending on the scalar of the sender or receiver node. Now, we can consider the hard attention block as a selector which for each node selects the best edge based on an ordered pair of ‘states priority’ (attention weights described above, which depend only on states of the corresponding nodes and edges) and $s_{ij}$. We note that this selector is related to the theoretical primitive $select\_best$ from RASP. We implement this simply by augmenting key vectors $K_{ij}$ of each edge with the indicator if the given edge has the “best” (min or max) scalar among the other edges to node $j$. Thus, scalars affect only the attention weights, not the messages and the node states.
@@
Given that, the interface of the proposed processor can be described as
@@
where $X^{t},X^{t+1}$ and $E^{t},E^{t+1}$ are from the fixed sets. State sets are independent of the execution step $t$ and the input graph (including scalar inputs $S$).
@@ 3.4 Manipulations over continuous inputs
The proposed selector offers a read-only interface to scalar inputs, which is not expressive enough for most of the algorithms. However, we note that the algorithms can be described as discrete manipulations over input data. For example, the Dijkstra algorithm (Dijkstra, 1959) takes edge weights as inputs and uses them to find the shortest path distances. Computed distances can affect the consequent execution steps. We note that such distances can be described as the sum of the weights of the edges that form the shortest path to the given node. In other words, the produced scalars depend only on input scalars and discrete execution states.
@@
In our experiments, we use a scalar updater capable of incrementing, moving, and adding scalars depending on discrete node/edge states:
@@
$s_{ij}^{t+1}$ $=\mathrm{inc}(e_{ij}^{t})+\mathrm{keep}(e_{ij}^{t})\cdot s_{ij}^{t}+\mathrm{push}(x_{i}^{t})\cdot s_{i}^{t},$
where $s_{i}$ are node-related scalars, $s_{ij}$ are edge-related scalars, and $\mathrm{inc}$, $\mathrm{keep}$, $\mathrm{push}$ are 0-1 functions representing if scalar in each node/edge should be incremented, kept, or pushed to any of its neighbors. We implement these functions as simple linear projections of node/edge features with consecutive discretization.
@@
$S^{t+1}=\mathrm{ScalarUpdate}(X^{t+1},E^{t+1},A,S^{t}).$
The proposed neural reasoners are fully discrete and can be interpreted by design. Moreover, the proposed selector block guarantees the predicted behavior of the message passing for any graph size, as it compares discrete state importance and uses continuous scalars only to break ties between equally important states.

### v3

@@ Abstract
Neural algorithmic reasoning aims to capture computations with neural networks by training models to imitate the execution of classical algorithms. While common architectures are expressive enough to contain the correct model in the weight space, current neural reasoners struggle to generalize well on out-of-distribution data. On the other hand, classical computations are not affected by distributional shifts as they can be described as transitions between discrete computational states. In this work, we propose to force neural reasoners to maintain the execution trajectory as a combination of finite predefined states. To achieve this, we separate discrete and continuous data flows and describe the interaction between them. Trained with supervision on the algorithm’s state transitions, such models are able to perfectly align with the original algorithm. To show this, we evaluate our approach on multiple algorithmic problems and achieve perfect test scores both in single-task and multitask setups. Moreover, the proposed architectural choice allows us to prove the correctness of the learned algorithms for any test data.
Machine Learning, ICML
@@ 1 Introduction
Learning to capture algorithmic dependencies in data and to perform algorithmic-like computations with neural networks is a core problem in machine learning, long studied using various approaches (Khardon & Roth, 1997; Graves et al., 2014; Zaremba & Sutskever, 2014; Reed & De Freitas, 2015; Kaiser & Sutskever, 2015; Veličković et al., 2020b).
Neural algorithmic reasoning (Veličković & Blundell, 2021) is a research area focusing on building models capable of executing classical algorithms. Relying on strong theoretical guarantees of algorithms to work correctly on any input of any size and distribution, this setting provides unlimited challenges for out-of-distribution generalization of neural networks. Prior work explored this setup using the CLRS-30 benchmark (Veličković et al., 2022), which covers classical algorithms from the Introduction to Algorithms textbook (Cormen et al., 2009) and uses graphs as a universal tool to encode data of various types. Importantly, CLRS-30 also provides the decomposition of classical algorithms into subroutines and simple transitions between consecutive execution steps, called hints, which can be used during training in various forms.
The core idea of the CLRS-30 benchmark is to understand how neural reasoners generalize well beyond the training distribution, namely on larger graphs. Classical algorithms possess strong generalization due to the guarantee that correct execution steps never encounter ‘out-of-distribution’ states, as all state transitions are predefined by the algorithm. In contrast, when encountering inputs from distributions that significantly differ from the training data, neural networks are usually not capable of robustly maintaining internal calculations in the desired domain. Consequently, due to the complexity and diversity of all possible data that neural reasoners can be tested on, the generalization performance of such models may vary depending on particular test distributions (Mahdavi et al., 2023). Given that, it is becoming important to interpret internal computations of neural reasoners to find errors or to prove the correctness of the learned algorithms (Georgiev et al., 2022).
Interpretation methods have been actively developing recently due to various real-world applications of neural networks and the need to debug and maintain systems based on them. Especially, the Transformer architecture (Vaswani et al., 2017) demonstrates state-of-the-art performance in natural language processing and other modalities, representing a field for the development of interpretability methods (Elhage et al., 2021; Weiss et al., 2021; Zhou et al., 2024; Lindner et al., 2023). Based on active research on a computational model behind the transformer architecture, recent works propose a way to learn models that are fully interpretable by design (Friedman et al., 2023).
@@
In this paper, we propose to force neural reasoners to follow the execution trajectory as a combination of finite predefined states, which is important for both generalization ability and interpretability of neural reasoners. To achieve that, we start with an attention-based neural network and describe three building blocks to enhance its generalization abilities: feature discretization, hard attention, and the separation of discrete and continuous data flows. In short, all mentioned blocks are connected:
@@
  • Hard attention is needed to ensure that attention weights will not be annealed for larger graphs and to limit the set of possible messages that each node can receive;
@@
Then, we build fully discrete neural reasoners for different algorithmic tasks and demonstrate their ability to perfectly mimic ground-truth algorithm execution. As a result, we achieve perfect test scores on multiple algorithmic tasks with guarantees of correctness on any test data. Moreover, we demonstrate that a single network is capable of executing all covered algorithms in a multitask setting, also achieving perfect generalization.
In summary, we consider the proposed blocks crucial components for robust and interpretable neural reasoners and demonstrate that discretized models, trained with hint supervision, perfectly capture the dynamics of the underlying algorithms and do not suffer from distributional shifts.
@@ 2.1 Algorithmic Reasoning
Performing algorithmic-like computations usually requires the execution of sequential steps and the number of such steps depends on the input size. To imitate such computations, neural networks are expected to be based on some form of recurrent unit, which can be applied to a particular problem instance several times (Zaremba & Sutskever, 2014; Kaiser & Sutskever, 2015; Vinyals et al., 2015; Veličković et al., 2020b).
@@
The CLRS-30 benchmark covers different algorithms over various domains (arrays, strings, graphs) and formulates them as algorithms over graphs. Additionally, CLRS-30 proposes utilizing the decomposition of the algorithmic trajectory execution into simple logical steps, called hints. Using this decomposition is expected to better align the model to desired computations and prevent it from utilizing hidden non-generalizable dependencies of a particular train set. Prior work demonstrates a wide variety of additional inductive biases for models towards generalizing computations, including different forms of hint usage (Veličković et al., 2022; Bevilacqua et al., 2023), biases from standard data structures (Jain et al., 2023; Jürß et al., 2024), knowledge transfer and multitasking (Xhonneux et al., 2021; Ibarz et al., 2022; Numeroso et al., 2023). Also, recent studies demonstrate several benefits of learning neural reasoners end-to-end without any hints at all (Mahdavi et al., 2023; Rodionov & Prokhorenkova, 2023).
The recently proposed SALSA-CLRS benchmark (Minder et al., 2023) enables a more thorough out-of-distribution (OOD) evaluation compared to CLRS-30 with increased test sizes (up to 100-fold train-to-test scaling, compared to 4-fold for CLRS-30) and diverse test distributions.
Despite significant gains in the performance of neural reasoners in recent work, current models still struggle to generalize to OOD test data (Mahdavi et al., 2023; Georgiev et al., 2023; Minder et al., 2023). While de Luca & Fountoulakis (2024) prove by construction the ability of the transformer-based neural reasoners to perfectly simulate graph algorithms (with minor limitations arising from the finite precision), it is still unclear if generalizable and interpretable models can be obtained via learning. Importantly, the issues of OOD generalization are induced not only by the challenges of capturing the algorithmic dependencies in the data but also by the need to carefully operate with continuous inputs. For example, investigating the simplest scenario of learning to emulate the addition of real numbers, Klindt (2023) demonstrates the failure of some models to exactly imitate the desired computations due to the nature of gradient-based optimization. This limitation can significantly affect the performance of neural reasoners on adversarial examples and larger input instances when small errors can be accumulated.
@@ 2.2 Transformer Interpretability and Computation Model
RASP (Weiss et al., 2021) is a programming language proposed as a high-level formalization of the computational model behind transformers. The main primitives of RASP are elementwise sequence functions, selectselect and aggregateaggregate operations, which conceptually relate to computations performed by different blocks of the model. Later, Lindner et al. (2023) presented Tracr, a compiler for converting RASP programs to the weights of the transformer model, which can be useful for evaluating interpretability methods.
While RASP might have limited expressibility, it supports arbitrarily complex continuous functions that in theory can be represented by the transformer architecture, but are difficult to learn. Also, RASP is designed to formalize computations over sequences of fixed length. Motivated by that, Zhou et al. (2024) proposed RASP-L, a restricted version of RASP, which aims to formalize the computations that are easy to learn with transformers in a size-generalized way. The authors also conjecture that the length-generalization of transformers on algorithmic problems is related to the ‘simplicity’ of solving these problems in the RASP-L language.
@@ 3.1 Encode-Process-Decode Paradigm
If the model is trained with hint supervision, the changes of node and edge features at each step are expected to be related to the original algorithm execution. In this sense, the processor network aims to mimic the algorithm’s execution in the latent space.
@@ 3.2 Discrete Neural Algorithmic Reasoners
Each node feature vector $x_{i}$ is projected into query $(Q_{i})$, key $(K_{i})$, and value $(V_{i})$ vectors via learnable parameter matrices $W_{Q}$, $W_{K}$, and $W_{V}$, respectively. Edge features $e_{ij}$ are projected into a key $(K_{ij})$ vector with a matrix $W_{K}^{E}$. Then, for each directed edge from a node $j$ to a node $i$ in the graph $G$, we compute the attention coefficient
@@
where $\langle a,b\rangle$ denotes the dot product. Then, each node $i$ normalizes all attention coefficients across its neighbors using the softmax function with temperature $\tau$ and receives the aggregated message:
$\hat{\alpha}_{ij}=\frac{\exp(\alpha_{ij}/\tau)}{\sum_{k\in\textit{N}(i)}\exp(\alpha_{ik}/\tau)},\quad M_{i}=\sum_{k\in\mathcal{N}(i)}\hat{\alpha}_{ik}V_{k},$ (1)
where $\mathcal{N}(i)$ denotes the set of all incoming neighbors of the node $i$ and $M_{i}$ is the message sent to the $i$-th node.
@@
Similar to Transformer Programs, we enforce attention to be hard attention. We found this property important not only for interpretability but also for size generalization, as hard attention allows us to overcome the annealing of the attention weights for arbitrarily large graphs and strictly limit the set of messages that each node can receive.
After the message computation, node and edge features are updated depending on the current values and received messages using feed-forward MLP blocks:
@@
We also enforce all node and edge features to be from a fixed finite set, which we call states. We ensure this property by adding discrete bottlenecks at the end of the processor block:
@@
We implement discretization by projecting the features into vectors of size $k$ (states logits) which we then use to produce a one-hot vector corresponding to a discrete state. During training with hints, we use teacher forcing and directly use ground-truth states as inputs for the next processor step. For no-hint learning, we apply Gumbel-Softmax (Jang et al., 2017) with annealing temperature to the states logits. During inference, we simply use the argmax function over the states logits dimension.
@@ 3.3 Continuous Inputs
Clearly, most algorithmic problems operate with continuous or unbounded inputs (e.g., weights on edges). Usually, all input data is encoded into node and edge features and the processor operates over the resulting vectors. The proposed discretization of such vectors would lead to the loss of information necessary for performing correct execution steps. To handle such inputs (which we will call scalars referring to both continuous and size-dependent integer inputs such as node indexes) one option is to use Neural Execution Engines (Yan et al., 2020) that enable operating with bit-wise representations of integer and (in theory) real numbers. Such representations are bounded by design, but fully discrete and interpretable.
We propose another option: to maintain scalar inputs (denote them by $S$) separately from the node and edge features and use them only as edge priorities $s_{ij}$ in the attention block. If scalars are related to the nodes, we assign them to edges depending on the scalar of the sender or receiver node. We implement this simply by augmenting the key vectors $K_{ij}$ of each edge with the indicator of whether the given edge has the “best” (min or max) scalar among the other edges to node $j$ with the same discrete state. Thus, scalars affect only the attention weights, not the messages or the node states.
@@
Thus, the interface of the proposed processor can be described as
@@
where $X^{t},X^{t+1}$ and $E^{t},E^{t+1}$ are from fixed sets. State sets are independent of the execution step $t$ and the input graph (including scalar inputs $S$).
@@ 3.4 Manipulations over Continuous Inputs
The proposed selector offers a read-only interface to scalar inputs, which is not expressive enough for most algorithms. However, we note that the algorithms can be described as discrete manipulations over input data. For example, the Dijkstra algorithm (Dijkstra, 1959) takes edge weights as inputs and uses them to find the shortest path distances. Computed distances can affect the subsequent execution steps. We note that such distances can be described as the sum of the weights of the edges that form the shortest path to the given node. In other words, the produced scalars depend only on input scalars and discrete execution states.
@@
In our experiments, we use a scalar updater capable of incrementing, moving, and adding scalars depending on discrete node/edge states: [^1] We additionally investigate if the operation set can be extended, see Appendix C.
@@
$s_{ij}^{t+1}$ $=\mathrm{inc}(e_{ij}^{t})+\mathrm{keep}(e_{ij}^{t})\cdot s_{ij}^{t}+\mathrm{push}(x_{i}^{t})\cdot s_{i}^{t},$ (2)
where $s_{i}$ are node-related scalars, $s_{ij}$ are edge-related scalars, and $\mathrm{inc}$, $\mathrm{keep}$, $\mathrm{push}$ are 0-1 functions representing whether a scalar in each node/edge should be incremented, kept, or pushed to any of its neighbors. We implement these functions as simple linear projections of node/edge features with subsequent discretization.
@@
$S^{t+1}=\mathrm{ScalarUpdate}(X^{t},E^{t},A,S^{t}).$
The proposed neural reasoners are fully discrete and can be interpreted by design. Moreover, the proposed selector block guarantees predictable behavior of the message passing across all graph sizes, as it compares discrete state importance and uses continuous scalars only to break ties between equally important states.